OpenStudy (anonymous):

Find an equation for the nth term of the sequence. -3, -12, -48, -192, ...

OpenStudy (anonymous):

@ganeshie8 @aaronq @Opcode

zepdrix (zepdrix):

So what is the common ratio? Looks like each term is the previous term multiplied by some value.

zepdrix (zepdrix):

\[\Large\rm r=\frac{a_2}{a_1}\] \[\Large\rm r=?\]

OpenStudy (mathmath333):

\(\large\rm its~~ geometric~~ progression~~ with~~ common~~ ratio(r) \\ \rm\large r=4~~ and ~~its ~~first ~~term(a)~~ is~~ -3 \) \(\rm \large ~~u ~~can~~ use ~~the ~~formula \) \(\rm\large~~ nth~~ term= ~~ar^{(n-1)}\)

OpenStudy (mathmath333):

\(\rm\large r=\dfrac{-12}{-3}=\dfrac{-48}{-12}.....\)